Case study: Parkinson’s Disease A 65-year-old patient with a 2-year history of PD has been experiencing an increase in symptoms, especially muscle tremors and rigidity. The patient has been taking carbidopa–levodopa (Sinemet CR) and selegiline since diagnosis. The physician has chosen to add benztropine to the patient’s medication regimen. The nurse is administering the first dose of benztropine and intends to provide patient teaching about the drug and its adverse effects. OTHER PERTINENT INFORMATION: Temperature - 98.6 degrees Fahrenheit Weight - 80 kg Height - 6 feet & 1 inch tall 1. What adverse effects are associated with benztropine? 2. Given the effect, the medication has on the pupils, what condition if present in the patient would cause concern? 3. An indication that the addition of benztropine is effective in this patient would be relief of which PD symptom? 4. Doctors order: 400mg of Levodopa OD PO and you only have 100mg tablets of Levodopa on hand. How many tablets would you administer to the patient? 5. Convert temperature into Celsius, his weight into pounds, and his height into centimeters. 6. What is the best pharmacological intervention to treat dyskinesias?
Answer #1

1.The adverse effects of benztropine are

  • Gastrointestinal :Nauseamdry mouth,constipation ,anorexia ,dysphagia
  • CNS:Delirium,cognitive changes,hallucination ,anxiety
  • Cardiac:increased pulse
  • Genitourinary : Urinary retention ,decreased sexual ability.
  • Sensory:eye pain,blurry vision
  • Severe :rash,itching,swelling,breathing difficulties ,abdominal pain

2.This medicine affects the vision of eye it causes blurry vision (patient can see rainbow colours at night)

It concerns If the patient has any eye problems like cataract which can lead to loss of vision

3.The patient will be relieved with rigidity and stiffness and walking is enabled.The major symp toms like tremors and bradykinesia will be relieved

4.Given:

Medication :Levodopa

Dose 400mg

Frequency :OD

Route:Oral

Available:100mg

If 1 tablet = 100mg

Then??=400mg

= 4 tablets (400÷100)

Ans:4 tablets has to be given

5.Temperature :98.6°F

C=F_32×5/9

=98.6_32×5/9

=66.6×5/9

=333÷9

=37°C

Weight =80kg

1kg=2.205 pounds

Multiply the weight by 2.205

=2.205×80

=176.37 pounds

Height= 6'1'

1 foot =30.48cm (30.48×6=182.88)

1 inch=2.54cm (2.54×1=2.54)

182.88+2.54

=185.42 cm

6.Extended release amantadine is found to be the best pharmacological intervention to treat dyskinesia.